I need help with solving linear systems. I need to use the linear combinations to solve the system of linear equations. I have tried several and they are not coming out right.
3a+9b=8b-a
5a-10b=4a-9b+5
Please help me!

Answer by oberobic(2304)

3a + 9b = 8b - a
5a - 10b = 4a - 9b + 5
.
for each equation, collect like terms and simplify to get into the right format
.
3a + 9b = 8b - a
add 'a' to both sides
4a + 9b = 8b
subtract 8b from both sides
4a + b = 0
.
5a - 10b = 4a - 9b + 5
subtract 4a from both sides
a - 10b = -9b + 5
add 9b to both sides
a - b = 5
.
now both equations are in the right form
4a + b = 0
a - b = 5
.
add them
5a = 5
a = 1
.
substitute back
.
4a + b = 0
4(1) + b = 0
4 + b = 0
b = -4
.
continue checking...
3a + 9b = 8b - a ??
3(1) + 9(-4) = 8(-4) - 1 ??
3 -36 = -32 -1 ??
-33 = -33 YES!
.
5a - 10b = 4a - 9b + 5 ??
5(1) - 10(-4) = 4(1) - 9(-4) + 5 ??
5 + 40 = 4 + 36 + 5 ??
45 = 45 YES!
.
Answer:
a = 1
b = -4